Farmer Fred said to Farmer John: “If you sell me 45 hectares of land, I will
have twice as much land as you.” Then Farmer John said to Framer Fred: “If
you sell me 45 hectares of land, I will have just as much land as you.” How
many hectares of land does farmer Fred have?
(A)135 (B) 180 (C) 225 (D) 270 (E) 315


Answer by ankor@dixie-net.com(22740)

Fred said to Farmer John: “If you sell me 45 hectares of land, I will have twice as much land as you.”
Then Farmer John said to Farmer Fred: “If you sell me 45 hectares of land, I will have just as much land as you.”
How many hectares of land does farmer Fred have?
:
Let f = amt of land Fred has
Let j = amt of land John has
:
“If you sell me 45 hectares of land, I will have twice as much land as you.”
f + 45 = 2(j - 45)
f + 45 = 2j - 90
f = 2j - 90 - 45
f = 2j - 135
:
John said to Framer Fred: “If you sell me 45 hectares of land, I will have just as much land as you.”
j + 45 = f - 45
j = f - 45 - 45
j = f - 90
:
Substitute (f-90) for j in the 1st equation
f = 2j - 135
f = 2(f-90) - 135
f = 2f - 180 - 135
f - 2f = -315
-f = - 315
f = 315 hectors Fred has